OpenStudy (anonymous):

WILL GIVE MEDAL IF CORRECT. I need help on two but I will fan and medal. What is the least common denominator of the fraction 2 over x plus the fraction x over x plus 1 plus the fraction x plus 1 over 3 x

OpenStudy (anonymous):

3x + 1 3(x2 + 1) 3(x + 3) 3x(x + 1)

OpenStudy (anonymous):

and After the fraction x plus 2 all over 3 minus the fraction x plus 1 all over 2x have been combined using the least common denominator of 6x, what is the numerator?

OpenStudy (anonymous):

2x2 + x − 3 2x2 + x − 1 −2x2 + x + 6 −x2 + 2x + 6

OpenStudy (owlcoffee):

Have you tried solving them, can you show me your work?

OpenStudy (anonymous):

I think it is c and a but I have no idea and I cant show my work because I am at an old computer right now:(

OpenStudy (owlcoffee):

\[\frac{ 2 }{ x }+\frac{ x }{ x+1 }+\frac{ x+1 }{ 3x }\] The "mcm" of this whole sum are all the not-repeated expressions, those being "x" , "3" and "x+1", therefore: \[mcm=(x)(x+1)(3)\] All you have to do is operate that succesive multiplication and you'll obtain the least common factor of the sum of all those fractions.

OpenStudy (anonymous):

Is it 3x+1?

OpenStudy (owlcoffee):

That's incorrect, note that there is an "x" multiplying the expression (x+1).

OpenStudy (anonymous):

3x(x + 1)?

OpenStudy (owlcoffee):

Correct.

OpenStudy (anonymous):

Thank you! Do you know the other one?

OpenStudy (owlcoffee):

Yes, but the question is if you know it.

OpenStudy (owlcoffee):

@misty1212 can explain it to you in simple terms.

OpenStudy (anonymous):

Would it be −2x2 + x + 6?

OpenStudy (owlcoffee):

Could you show me how'd you work it out?